# SimuCourt: Building Judicial Decision-Making Agents with Real-world Judgement Documents

### v1

@@ Abstract
With the development of deep learning, natural language processing technology has effectively improved the efficiency of various aspects of the traditional judicial industry. However, most current efforts focus solely on individual judicial stage, overlooking cross-stage collaboration. As the autonomous agents powered by large language models are becoming increasingly smart and able to make complex decisions in real-world settings, offering new insights for judicial intelligence. In this paper, (1) we introduce SimuCourt, a judicial benchmark that encompasses 420 judgment documents from real-world, spanning the three most common types of judicial cases, and a novel task Judicial Decision-Making to evaluate the judicial analysis and decision-making power of agents. To support this task, we construct a large-scale judicial knowledge base, JudicialKB, with multiple legal knowledge. (2) we propose a novel multi-agent framework, AgentsCourt. Our framework follows the real-world classic court trial process, consisting of court debate simulation, legal information retrieval and judgement refinement to simulate the decision-making of judge. (3) we perform extensive experiments, the results demonstrate that, our framework outperforms the existing advanced methods in various aspects, especially in generating legal grounds, where our model achieves significant improvements of 8.6% and 9.1% F1 score in the first and second instance settings, respectively.
SimuCourt: Building Judicial Decision-Making Agents with Real-world Judgement Documents
[…]
@@ 1 Introduction
Figure 1: We formulate the Judicial Decision-Making task using the real-world judgement documents: given the case details above, judge agent must 1) conduct a logically clear case analysis; 2) provide precise legal grounds; 3) issue a definitive judgement.
@@
Recent advances in deep learning have significantly impacted the legal domain, with notable achievements in legal event detection Yao et al. (2022a); Fei et al. (2023), legal question answering Zhong et al. (2020); Khazaeli et al. (2021); Martinez-Gil (2023), and legal judgment prediction Chalkidis et al. (2019); Hwang et al. (2022); Fei et al. (2023). These developments have effectively alleviated the long-standing issue in the judicial industry of "too many cases, too few legal professionals". However, case trial is a coherent process involving multiple stages such as court debates, case analysis, and precedents retrieval. The complexity of this process demands close collaboration and interaction between stages. Although current research has made progress in individual areas, it often overlooks the inherent connections between these stages of the trial process. This results in the need to rely on the deep involvement of legal experts when dealing with complex judicial decisions. Meanwhile, autonomous agents based on large language models (LLMs) have shown considerable progress in various traditional natural language processing (NLP) tasks Brown et al. (2020); Wei et al. (2022); Wang et al. (2023); Qian et al. (2023); Wu et al. (2023). An increasing number of agents are being proposed to make decisions in real-world environments Yao et al. (2023); Richards (2023); Chen et al. (2023), which offers new insights for judicial intelligence.
@@
However, simulating judicial decision-making is a non-trivial task because agents must navigate complex situations involving multiple stakeholders, understand the subtle nuances of legal provisions, and consider ethical and social justice factors. This presents three unique challenges to the agent system: (1) Expert knowledge of judicial domain. Judicial adjudication requires an in-depth understanding and accurate application of specialized knowledge such as laws, case precedents, and judicial procedures. (2) Complex and hybrid reasoning. The agents must be capable of handling a complex amalgamation of logical, factual, and legal reasoning, often interwoven in cases. (3) Intricate ethical relationships. In judicial decision, ethical and moral considerations, which are often subtle and multi-faceted, must be taken into account.
In this paper, we introduce SimuCourt, a judicial benchmark designed to evaluate Agent-as-Judge across a spectrum of different cases. SimuCourt encompasses 420 judgement documents, spanning the three most common types of judicial cases — criminal, civil, and administrative — in both first-instance and second-instance (appellate) courts, as well as covering three key societal roles: government agencies, the prosecutor’s office, and individuals. Specifically, criminal cases deal with acts that violate laws. Civil cases typically involve disputes between individuals, such as contract disputes or torts. Administrative cases concern disputes between individuals and government agencies. All the cases come from the China Judgements Online  [^1] https://wenshu.court.gov.cn/, which is an official platform established by the Supreme People’s Court of China, aimed at publicly releasing the judgement documents of courts at all levels in China. We formulate a novel task Judicial Decision-Making, as illustrated in Figure 1. Given the case details, agent must conduct a logically clear case analysis, provide precise legal grounds and issue a definitive judgement. Furthermore, we construct a large-scale judicial knowledge base, JudicialKB, to support this domain task. It encompasses a variety of legal knowledge, including effective laws and regulations, highly cited judicial papers, and precedents from recent years. The use of real data allows the agents developed on it can be transferred into real applications without any gaps.
@@
We also present a novel multi-agent framework, AgentsCourt. It follows the classic court trial process and simplifies the process into four phases: opening remarks, court debate, precedent retrieval, and judgement, as depicted in Figure 2. Specifically, we first develop a Court Debate Simulation Module with three agents. One agent serves as the judge to open a court session and announce the basic facts of the case. The other two agents are designed as the plaintiff and the defendant (or the-accused in criminal cases) respectively, and articulate their points of view during the court debate phase. This module provides a platform for all parties involved to present their points and arguments fairly. Then, we devise the Legal Information Retrieval Module which employs a judge assistant agent to integrate the most relevant precedents, articles and other information retrieved from the knowledge base we constructed and the internet. Next, we propose the Judgement Refinement Module which firstly makes a preliminary judgement according to the inherent judicial expertise of the agent elicited by the established facts of current case and the transcripts of court debate, then subsequently refines the judgement using legal information retrieved.
@@
  • We introduce SimuCourt, a judicial benchmark encompasses the three most common types of cases, enabling reliable assessment of the judicial analysis and decision-making power of agents for real judicial practice.
  • We propose a novel multi-agent framework AgentsCourt. Given the basic information of a case, AgentCourt can sequentially simulate court debate, retrieve precedents, analyze cases, provide legal grounds, and deliver clear judgment. The new judicial paradigm simplifies the process of making judicial decisions, significantly enhancing judicial efficiency.
  • We perform experiments and ablation studies to explore factors that impact performance. The results indicate that our framework outperforms the existing advanced methods in various aspects, especially in generating legal grounds, where our system achieves significant improvements of 8.6% and 9.1% F1 score in the first and second instance experimental settings, respectively. Dataset and codes will be available at https://github.com/Zhitao-He/SimuCourt.
@@ 2 Related Work
Real-world Agent tasks  Agents are becoming smarter and more autonomous, focusing on practical real-world task beyond traditional tasks. Shridhar et al. (2020) introduces a simulator that enables agents to learn abstract, text-based policies in TextWorldCôté et al. (2019) and then execute house-holding goals. Toyama et al. (2021) proposes the use of an Android simulator to learn phone operations. Yao et al. (2022b); Zhou et al. (2023); Deng et al. (2023) create simulated environments to develop web agents that perform web browsing tasks. Liu et al. (2023) examine agents’ abilities to operate on real databases via SQL and evaluate agents in genuine OS’ interactive bash environments. Cheng and Chin (2024) introduces a lifelong learning environment for developing autonomous agents capable of performing human-like analysis on societal topics such as economics.
@@
Multi-agent framework  Cooperation among agents like human group dynamics can enhance the efficiency and effectiveness of task accomplishment. Li et al. (2023) enables two communicative agents to engage in a conversation and cooperate with each other to solve assigned tasks. Park et al. (2023) found social behaviors autonomously emerge within a group of agents. Qian et al. (2023); Hong et al. (2023) present innovative paradigms that leverages LLMs throughout the entire software development process by natural language communication. Du et al. (2023); Zhang et al. (2023); He et al. (2023); Chen et al. (2023); Wu et al. (2023) further leverage multi-agent cooperation to achieve better performance on multiple tasks.
Agent system for judicial industry  The rapid emergence and growing popularity of language models have laid the groundwork for the development of specialized legal language models Cui et al. (2023); Nguyen (2023); Huang et al. (2023). Although most existing work primarily uses large models as unstructured knowledge bases for legal question answering and the results demonstrate that LLMs have surpassed 90% of human-level performance in the Uniform Bar Exam Achiam et al. (2023), there is a concern that information retrieval systems might select the correct answer for incorrect reasons Hubbard et al. (2017); Couch et al. (2018). Furthermore, complex judicial processes, such as judicial decision-making, are still mainly conducted by legal experts. In this paper, we focus on modeling the judicial decision-making process as an agent generative task, which involves analyzing case details, providing judicial grounds, and determining judgement.
@@ 3.1 Data Collection
We collect 420 real-world cases from the China Judgements Online, which span across two fundamental trial stages: first instance and second instance. These cases encompass three types: criminal, civil, and administrative. For first-instance cases, each sample includes the Complaint, Statement of Plaintiff and Defendant, Determine facts, etc. For second-instance cases, each sample contains Petition for appeal, Statement of the appellant and appellee, Determine facts in the first Instance, etc. Detailed list can be found in the Appendix A. Most of cases were released after April 2023. This minimizes the risk of data leakage. [^2] The cutoff date of pretraining data for gpt-3.5-turbo-0613 and gpt-4-1106-preview is officially before April 2023.
3.2 Data Description
Our choice of cases is driven by three reasons: (1) Diversity of causes of action. Based on our statistical analysis of data from the China Judgements Online over the past few years, we observed a significant long-tail distribution in various types of cases. For example, as shown in Appendix C, in the total civil cases of 2022, the top 15 causes of action accounted for 66% of the total number of cases. To reflect a broader spectrum of legal practice, we focus on maintaining diversity in the types of causes of action; (2) Clarity of case analysis and facts. We have meticulously selected judgement documents that provide detailed case analysis and clear determine facts for annotation. This aim is to enhance the quality and accuracy of data annotation while aiding agents in better understanding the judicial reasoning and legal grounds; (3) Uniqueness and accuracy of judgements. We prioritize cases that are not overturned in appellate review. This ensures the consistency of our evaluation, as these cases have already undergone a rigorous litigation process and the judgements are fair. Detailed data statistics of SimuCourt are shown in Table 1.
3.3 Data Quality
Our selected judgement documents undergo rigorous scrutiny, ensuring the accuracy and completeness of the legal texts and information. The clarity of information in these documents facilitates our efficient and precise data annotation. We first process the privacy information of all documents. Specifically, We have meticulously anonymized sensitive information in the judgement documents. Then, After completing data annotation and handling private information, we manually inspect the quality of SimuCourt from various aspects. Detailed data quality inspection can be found in Appendix B.
@@ 3.3 Data Quality
Avg # of Legal grounds 6.3 3.3 1.6
Max # of Legal grounds 11 10 8
Total # of Legal grounds 198 153 92
@@
Table 1: Statistics of SimuCourt.
Figure 3: Overview of our multi-agent framework. The Court Debate Simulation Module recreates the court debate process through role-playing, mining different parties’ points from limited real records. The Legal Information Retrieval Module employ an assistant agent to integrate information retrieved. The Judgement Refinement Module exploit the inherent judicial expertise of the judge agent and refines the judgment using information retrieved.
3.4 Judicial Knowledge Base Construction
To make accurate judicial decisions, judges must possess extensive legal knowledge. Furthermore, given the diversity and complexity of human society, each case may involve different facts, parties, and locations. To this end, we construct a large scale judicial knowledge base consists of laws, regulations, judicial interpretation, journal articles, and precedents. Detailed data statistics of JudicialKB are shown in Table 2.

### v3

# AgentsCourt: Building Judicial Decision-Making Agents with Court Debate Simulation and Legal Knowledge Augmentation

@@ Abstract
With the development of deep learning, natural language processing technology has effectively improved the efficiency of various aspects of the traditional judicial industry. However, most current efforts focus on tasks within individual judicial stages, making it difficult to handle complex tasks that span multiple stages. As the autonomous agents powered by large language models are becoming increasingly smart and able to make complex decisions in real-world settings, offering new insights for judicial intelligence. In this paper, (1) we propose a novel multi-agent framework, AgentsCourt, for judicial decision-making. Our framework follows the classic court trial process, consisting of court debate simulation, legal resources retrieval and decision-making refinement to simulate the decision-making of judge. (2) we introduce SimuCourt, a judicial benchmark that encompasses 420 Chinese judgment documents, spanning the three most common types of judicial cases. Furthermore, to support this task, we construct a large-scale legal knowledge base, Legal-KB, with multi-resource legal knowledge. (3) Extensive experiments show that our framework outperforms the existing advanced methods in various aspects, especially in generating legal articles, where our model achieves significant improvements of 8.6% and 9.1% F1 score in the first and second instance settings, respectively.
AgentsCourt: Building Judicial Decision-Making Agents with Court Debate Simulation and Legal Knowledge Augmentation
Zhitao He1,2, Pengfei Cao1,2, Chenhao Wang1,2, Zhuoran Jin1,2, Yubo Chen1,2 Jiexin Xu3 ,Huaijun Li3, Xiaojian Jiang3, Kang Liu1,2, Jun Zhao1,2 1 The Laboratory of Cognition and Decision Intelligence for Complex Systems, Institute of Automation, Chinese Academy of Sciences, Beijing, China 2 School of Artificial Intelligence, University of Chinese Academy of Sciences, Beijing, China 3 AI Lab, China Merchant Bank, ShenZhen, China {zhitao.he, pengfei.cao, yubo.chen, kliu, jzhao}@nlpr.ia.ac.cn
Figure 1: We formulate the Judicial Decision-Making task using the real-world judgement documents: given the case details above, judge agent must 1) conduct a logically clear case analysis; 2) provide precise legal articles; 3) issue a definitive judgement.
@@ 1 Introduction
Framework AgentsCourt LaWGPT PLJP HRN RLJP
(This work) (Song et al., 2023) (Wu et al., 2023b) (Lyu et al., 2023) (Wu et al., 2022)
Case Analysis
Precedent Retrieval
Web Research
Court Simulation
Judgement Prediction
Legal Articles Generation Multiple Single Single Single Single
Case Type Various Various Crime Crime Crime
Table 1: A comparison of our AgentsCourt to notable legal domain frameworks.
@@
Recent advances in deep learning have significantly impacted the legal domain, with notable achievements in legal question answering Zhong et al. (2020b); Khazaeli et al. (2021); Cui et al. (2023), legal case retrieval Sugathadasa et al. (2019); Shao et al. (2020); Li et al. (2023b); Shao et al. (2023) and legal judgment prediction Xiao et al. (2018); Chalkidis et al. (2019); Wu et al. (2022, 2023b). These developments have effectively alleviated the long-standing issue in the judicial industry of "too many cases, too few legal professionals". However, case trial is a coherent process involving multiple stages such as court debates, case analysis, and legal judgment prediction. The complexity of this process demands close collaboration and interaction between stages. Although current research has made progress in individual stages, it often overlooks the inherent connections between these stages of the trial process. This results in the need to rely on the deep involvement of legal experts when dealing with complex judicial decisions. Meanwhile, autonomous agents based on large language models (LLMs) have shown considerable progress in various traditional natural language processing (NLP) tasks Brown et al. (2020); Wei et al. (2022); Wang et al. (2023); Qian et al. (2023); Wu et al. (2023a) and making decisions in real-world environments Yao et al. (2023); Richards (2023); Chen et al. (2023), which offers new insights for judicial intelligence.
@@
However, simulating judicial decision-making is a non-trivial task because agents must navigate complex situations involving multiple stakeholders, understand the subtle nuances of legal provisions, and consider ethical and social justice factors. This presents three unique challenges to the agent system: (1) Intricate ethical relationships. In judicial decision, ethical and moral considerations, which are often subtle and multi-faceted, must be taken into account.(2) Expert knowledge of judicial domain. Judicial adjudication requires an in-depth understanding and accurate application of specialized knowledge such as laws, regulations and precedents. (3) Complex and hybrid reasoning. The agents must be capable of handling a complex amalgamation of logical, factual, and legal reasoning, often interwoven in cases.
To tackle the aforementioned challenges, we propose a novel multi-agent framework, AgentsCourt, for the Judicial Decision-Making task. As illustrated in Figure 1, given the case details, the task requires the agent to conduct a logically clear case analysis, provide precise legal articles and issue a definitive judgement. AgentsCourt follows the classic court trial process: opening remarks, court debate, precedent retrieval, and judgement, as depicted in Figure 2. Specifically, we first develop a Court Debate Simulation Module with three agents, which serves as a platform for all parties involved to present their points to clarify the intricate ethical relationships in the case. One agent serves as the judge to open a court session and announce the basic facts of the case. The other two agents are designed as the plaintiff and the defendant respectively, and articulate their points of view during the court debate phase. Then, we devise the Legal Resources Retrieval Module to address the inadequacy of expert knowledge. This module employs a judge assistant agent to integrate the most relevant precedents, articles and other information retrieved from the knowledge base we constructed and the internet. Next, we propose the Decision-Making Refinement Module to facilitate complex and hybrid reasoning. This module firstly makes a preliminary judgement according to the inherent judicial expertise of the agent elicited by the established facts of current case and the transcripts of court debate, then subsequently refines the judgement using legal information retrieved.
The comparison between our framework and prior works is listed in Table 1. It is worth noting that our framework is not tailored to a specific legal system. AgentsCourt can achieve court simulation, precedent retrieval, judgment prediction, and supports the generation of multiple legal articles for practical judicial practice.
@@
We also introduce SimuCourt, a judicial benchmark designed to evaluate Agent-as-Judge across a spectrum of different cases. SimuCourt encompasses 420 Chinese judgement documents, spanning the three most common types of judicial cases — criminal, civil, and administrative — in both first-instance and second-instance (appellate) courts, as well as covering three key societal roles: government agencies, the prosecutor’s office, and individuals. Specifically, criminal cases involve acts that are identified as violations of criminal law, such as theft. Civil cases typically involve disputes between individuals, such as contract disputes or torts. Administrative cases concern disputes between individuals and government agencies. All the cases come from the China Judgements Online  [^1] https://wenshu.court.gov.cn/, which is an official platform established by the Supreme People’s Court of China, aimed at publicly releasing the judgement documents of courts at all levels in China. Furthermore, we construct a large-scale legal knowledge base, Legal-KB, to support this domain task. It encompasses a variety of legal knowledge, including effective laws and regulations, highly cited judicial papers, and precedents from recent years. The use of real data allows the agents developed on it can be transferred into real applications without any gaps.
@@
  • We propose a novel multi-agent framework AgentsCourt. Given the basic information of a case, our framework can sequentially simulate court debate, retrieve precedents, analyze cases, provide legal articles, and deliver clear judgment. The new judicial paradigm simplifies the process of making judicial decisions, significantly enhancing judicial efficiency.
  • We introduce SimuCourt, a judicial benchmark encompasses the three most common types of cases, enabling reliable assessment of the judicial analysis and decision-making power of agents for real judicial practice. Furthermore, we construct a legal knowledge base, Legal-KB, with multi-resource legal knowledge to support this task.
  • We perform extensive experiments and ablation studies. The results indicate that our framework outperforms the existing advanced methods in various aspects, especially in generating legal articles, where our system achieves notable improvements of 8.6% and 9.1% F1 score in the first and second instance experimental settings, respectively.
@@ 2 Related Work
Legal Artificial Intelligence  Legal Artificial Intelligence seeks to improve legal tasks by employing artificial intelligence techniques Surden (2019); Zhong et al. (2020a); Katz et al. (2023). With the continuous development of deep learning, the legal field has witnessed the emergence of more intelligent applications across various legal tasks. These tasks span across areas such as legal judgment prediction (LJP) Xiao et al. (2018); Zhong et al. (2018); Xu et al. (2020); Yue et al. (2021); Wu et al. (2022, 2023b), legal question answering Zhong et al. (2020b); Cui et al. (2023); Louis et al. (2024); Fei et al. (2023), legal language understanding Chalkidis et al. (2022); Xiao et al. (2021); Niklaus et al. (2023); Yu et al. (2023), legal case retrieval Sugathadasa et al. (2019); Shao et al. (2020); Li et al. (2023b); Shao et al. (2023), legal document summarization Kanapala et al. (2019); Jain et al. (2023, 2024). While these existing efforts have made progress in individual legal tasks, they have overlooked the interconnection between different tasks, resulting in the necessity to heavily rely on the deep involvement of legal experts when dealing with complex judicial decisions. In this work, we focus on completing the entire process of judicial decision-making through multi-agent collaboration.
@@
Multi-agent framework  Cooperation among agents like human group dynamics can enhance the efficiency and effectiveness of task accomplishment. Li et al. (2023a) enables two communicative agents to engage in a conversation and cooperate with each other to solve assigned tasks. Park et al. (2023) found social behaviors autonomously emerge within a group of agents. Qian et al. (2023); Hong et al. (2023) present innovative paradigms that leverages LLMs throughout the entire software development process by natural language communication. Du et al. (2023); Zhang et al. (2023); He et al. (2023); Chen et al. (2023); Wu et al. (2023a) further leverage multi-agent cooperation to achieve better performance on multiple tasks.
2.1 Task Formulation
We propose a generative task to evaluate agent as judge. Specifically, as shown in Figure 1, we formulate the Judicial Decision-Making task as given the case details of a case, such as Determine facts, Complaint/Indictment, Statement of the plaintiff and the defendant, the agent system needs to make a complete judicial decision, which includes a clear and reasonable case analysis, rigorous legal articles, and definitive final judgement. SimuCourt encompasses two experimental settings:
Figure 3: Overview of our multi-agent framework. The Court Debate Simulation Module recreates the court debate process through role-playing, mining different parties’ points from limited real records. TheLegal Resources Retrieval Module employ an assistant agent to integrate information retrieved. The Decision-Making Refinement Module exploit the inherent judicial expertise of the judge agent and refines the judgment using information retrieved.
First Instance This setting refers to the trial court level, where the judge determines the guilt of the defendant, and assesses whether punitive measures are warranted. Within this setting, the primary focus is on evaluating the agent’s understanding and analysis of case facts.
Second Instance This setting refers to the appellate court level. During this stage, the judge re-evaluates the case, considering new evidence. The objective at this stage is to ensure the legality and fairness of the initial judgement, identifying legal errors or inappropriate application of regulations from the first instance and demonstrating the capability to effectively handle new evidence.
3 The AgentsCourt Framework
3.1 Court Debate Simulation
The court debate provides a platform for all parties involved to present their points and arguments comprehensively and fairly, which can significantly influence the judgement of the case.
Court Simulation Due to the majority of judgement documents only recording the key points of the plaintiff’s and defendant’s statements, obtaining complete court transcripts is challenging. Fortunately, as large language models have shown remarkable ability in role-playing Li et al. (2023a); Qian et al. (2023); Chen et al. (2023), in this module, we aim to reconstruct the court debate with multiple agents for each case. We set up three agents to play the roles of the judge, plaintiff, and defendant respectively. For each agent, we carefully design an role-playing prompt to build their character personality and use the actual statements from judgment documents as the their starting prompts. It is worth noting that due to the limited record of statements in judgment documents, we combine the plaintiff and their representative, as well as the defendant and their representative, into the plaintiff and defendant, respectively, without setting separate roles for representatives.
Court Debate In this stage, both the plaintiff and the defendant need to present their arguments in line with their interests. The plaintiff should vigorously argue their complaint, articulating their stance and reasoning. Meanwhile, the defendant must defend their actions, aiming to prove their innocence or seek a lighter penalty. During the court session, the judge agent first delivers opening remarks, which include basic information about the plaintiff and the defendant, determination of facts, and so on. Then, the trial moves into the court debate stage and the communication between the agents will be recorded as court transcripts. We present an example of court transcripts in Table 12.
3.2 Legal Resources Retrieval
Court debate serves as a platform to thoroughly explore the facts and contentious issues within a case, making the judge better comprehend the complexity of the matter. Furthermore, to make accurate judicial decisions, judges must possess extensive legal knowledge and case information.
Figure 4: Automatic retrieval of precedents.
Automatic Information Retrieval In terms of knowledge base retrieval, as presented in Figure 4, the assistant first predict the type of case based on the determine facts of the current case. Due to the vast number of documents in the knowledge base, and the fact that cases with the same cause often have more similar keywords, we employs the BM25 model Lin et al. (2021) for efficient rough retrieval to obtain the top 100 documents from the knowledge base. Building on this, we further utilize the BGE-Large model Xiao et al. (2023) to encode and re-rank these retrieved documents and choose the most relevant document to the current case as the optimal precedent. Additionally, to obtain more comprehensive laws and regulations relevant to the current case without introducing additional context, the judge assistant extracts the corresponding legal articles from the top 5 precedents as related legal provisions of current case.
3.3 Judgement Refinement
Preliminary Judgement As shown in the bottom of Figure 3, after receiving the determine facts of current case and transcripts of simulated court debate, the judge agent takes the action of analysis, then provides its legal articles and subsequently reaching a preliminary judgement.
@@ 4 The SimuCourt Benchmarck
The task, Judicial Decision-Making, requires agents to conduct case analysis, generate legal articles and judgments. However, most existing legal datasets suffer from several limitations when it comes to assessing the Agent-as-Judge paradigm: 1) only contain the factual information of cases; 2) only focus on criminal cases; 3) only evaluate judgments. To this end, we propose SimuCourt, a judicial benchmark for a reliable assessment of the judicial analysis and decision-making power of agents. A comparison between our dataset and previous works is presented in Table 2.
@@ 4.1 Data Collection
We collect 420 real-world cases from the China Judgements Online, which span across two fundamental trial stages: first instance and second instance. These cases encompass three types: criminal, civil, and administrative. For first-instance cases, each sample includes the indictment, the point of the defendant, determine facts, etc. For second-instance cases, each sample contains petition for appeal, the point of the appellant and appellee, etc. Detailed list and data examples can be found in the Appendix C. Most of cases were released after April 2023. This minimizes the risk of data leakage [^2] The cutoff date of pretraining data for gpt-3.5-turbo-0613 and gpt-4-1106-preview is officially before April 2023.. Detailed data statistics of SimuCourt are shown in Table 3. Furthermore, our dataset undergo rigorous scrutiny, ensuring the accuracy and completeness of the legal texts and information. Details of data collection and quality inspection can be found in Appendix D.
Resource SimuCourt CAIL SLJA-SYN
Background of Defendant?
Statement of Different Parties?
Multi-article Scenario?
Case Analysis Evaluation?
Judgement Evaluation
Laws Involved? 443 1 1
Case Retrival? 6.5M 2.6M
Various Case Types? Crime, Civial, Admini. Crime Crime
Different Instances Involved? First/Second First First
Table 2: A comparison of our SimuCourt to remarkable legal domain datasets. CAIL Xiao et al. (2018) is a widely used legal judgment prediction dataset, where each case comes with a fact description; SLJA-SYN Deng et al. (2023) is a comprehensive legal dataset designed to support multiple tasks such as article retrieval, article interpretation generation, criminal element generation and legal judgment prediction.
@@
Avg # of Legal articles 6.3 3.3 1.6
Max # of Legal articles 11 10 8
Total # of Legal articles 198 153 92
@@
Table 3: Statistics of SimuCourt. Length is measured via the number of words
4.2 Legal Knowledge Base Construction
To make accurate judicial decisions, judges must possess extensive legal knowledge. Furthermore, given the diversity and complexity of human society, each case may involve different facts, parties, and locations. To this end, we construct a large scale legal knowledge base consists of laws, regulations, judicial interpretation, journal articles, and precedents. Detailed data statistics of Legal-KB are shown in Table 4.